August 26, 2012 by
I went for one last all-night observing session before the semester starts; the weather was great, no clouds and pretty good seeing. No issues with the telescope. The night was spent observing a few different variable stars.
I ended the night with some twilight flats and some floodlight flats.
August 21, 2012 by
Dr. Craig, Dr. Cabanela, Aaron and I went out to the observatory tonight. Aaron and I spent the first half of the night adjusting the set star, which was supposed to appear on the illuminated slit.We got to see the star, but it was out of focus. When we started changing the focus, we completely lost track of it.
August 16, 2012 by
Dr. Craig and I went to the observatory tonight. The sky was partially cloudy. Transparency was moderately good. We mounted the spectrograph and spent some time setting the focus. The focus was 3256 at an ambient temperature of 68 F. Again, always cross check the sky clock with FM weather website before you depart.
August 12, 2012 by
I got back from DC yesterday. Dr. Craig, Dr. Winkler and I went to the observatory tonight. We mounted the spectrograph and made some progress in collecting spectra of a set star (del Boo). The new connector tube lowered the focus around 3234.